Refer to the data set of times, in minunes required for an airplane to tadl out for takeoff, Isted below Find the mean and median How is it holphul to find the mean? Cick the kcon for the taxd out takeolt data Find the mean and median of the data set using a calculator or sinilar data analysis techimologyMore Ino The mean of the data set is minutes Round to one decimal place as needed) The median of the data set is minutes Round to one decimal place as needed) 34 23 37 40 20 15 16 30 31 35 19 14 13 13 15 30 23 1 36 23 28 16 23 13 50 30 33 37 42 31 19 25 37 22 7 37 How is it helpful to find the mean? 15 21 14 21 39 35 31 23 35 37 28 12 O A. The mean taxi out time is important for determining the quality of the flight 。B. The mean taxi out time is important for cakulating and shedding the arrival time O C. The mean taxi out time is important for the condition of the runway Print Done O D. The mean taxi out time is not important
Answer #1

Answer

First, we need to arrange the data in ascending order or increasing order.

11
12
13
13
13
14
14
15
15
15
16
16
19
19
20
21
21
22
23
23
23
23
23
25
28
28
30
30
30
31
31
31
33
34
35
35
35
36
37
37
37
37
37
39
40
42
47
50

We know that the formula for the mean is given as

Mean = (sum of all data values)/(total number of given values) = (11+12+...+47+50)/48 = 1279/48 = 26.6

So, Mean value is 26.6

Formula for median is given as

Median = ((n/2)th + ((n/2)+1)th)/2

where n is the sample size= 48

setting the given values, we get

Median = ((48/2)th (48/2) 1)th)/2 = (24th term + (24+1)th term)/2 = (24th term + 25th term)/2

where 24th term = 25 and 25th term = 28

setting the values, we get

Median = (25+28)/2 = 53/2 = 26.5

So, median is 26.5

Mean time is useful in the calculation and scheduling the arrival time because it will help passengers in saving their time.

So, option B is correct answer.